# Visual Generation Without Guidance

### v1

@@ Abstract
Classifier-Free Guidance (CFG) has been a default technique in various visual generative models, yet it requires inference from both conditional and unconditional models during sampling. We propose to build visual models that are free from guided sampling. The resulting algorithm, Guidance-Free Training (GFT), matches the performance of CFG while reducing sampling to a single model, halving the computational cost. Unlike previous distillation-based approaches that rely on pretrained CFG networks, GFT enables training directly from scratch. GFT is simple to implement. It retains the same maximum likelihood objective as CFG and differs mainly in the parameterization of conditional models. Implementing GFT requires only minimal modifications to existing codebases, as most design choices and hyperparameters are directly inherited from CFG. Our extensive experiments across five distinct visual models demonstrate the effectiveness and versatility of GFT. Across domains of diffusion, autoregressive, and masked-prediction modeling, GFT consistently achieves comparable or even lower FID scores, with similar diversity-fidelity trade-offs compared with CFG baselines, all while being guidance-free. Code will be available at https://github.com/thu-ml/GFT.
@@
Figure 1: Impact of adjusting sampling temperature $\beta$ for DiT-XL/2 after applying GFT. More examples are in Figure 11.
@@ 1 Introduction
Low-temperature sampling is a critical technique for enhancing generation quality by focusing only on the model’s high-likelihood areas. Visual models mainly achieved this through Classifier-Free Guidance (CFG) (Ho & Salimans, 2022). CFG jointly optimizes both conditional and unconditional models during training and combines them to define the sampling process. By altering the guidance scale, it can flexibly trade off image fidelity and diversity at inference time, while significantly improving sample quality. Due to its effectiveness, CFG has been a default technique for a wide spectrum of visual generative models, including diffusion (Ho et al., 2020), autoregressive (AR) (Chen et al., 2020; Tian et al., 2024), and masked-prediction models (Chang et al., 2022; Li et al., 2023).
Despite the wide application, CFG requires inferencing both conditional and unconditional models to achieve the sampling distribution. This not only doubles the sampling cost but also complicates the post-training of visual models: When distilling pretrained diffusion models (Meng et al., 2023; Luo et al., 2023; Yin et al., 2024b) for fast inference or applying RLHF techniques (Black et al., 2023; Chen et al., 2024b), CFG requires the extra unconditional model to be additionally considered in the algorithm design.
It’s also worth noting that for language models (LMs), a single model is sufficient to represent sampling distributions across various temperatures. However, similarly following LMs’ approach to divide model output by a constant temperature value has largely been found ineffective in visual sampling (Dhariwal & Nichol, 2021), even for visual AR models with similar architecture to LMs (Sun et al., 2024). This leaves us wondering, how can we control the sampling temperature for visual models using only one model?
@@
Existing attempts mainly include distillation methods for diffusion models (Meng et al., 2023; Luo et al., 2023; Yin et al., 2024b) and alignment methods for AR models (Chen et al., 2024b). These methods rely heavily on pretrained CFG networks for loss definition and do not support training guidance-free models from scratch. Their two-stage optimization pipeline may also lead to performance loss compared with CFG, even after extensive tuning. Generalizability is also a concern. Current methods are tailored for either continuous diffusion models or discrete AR models, lacking the versatility to cover all domains.
@@
We propose Guidance-Free Training (GFT) as a foundational algorithm for building guidance-free visual generative models. GFT matches CFG in performance while requiring only a single model for temperature-controlled sampling (Figure 1), effectively halving sampling costs. It offers stable and efficient training with the same convergence rate as CFG, almost no extra memory usage, and only 10–20% additional computation per training update. Notably, GFT is highly versatile, applicable in all visual domains within CFG’s scope, including diffusion, AR, and masked models.
@@
The core idea behind GFT is straightforward (Figure 2). GFT and CFG optimize the same conditional loss. CFG parameterizes an independent conditional model and combines it with another unconditional model for guided sampling. Contrastively, GFT does not explicitly define a conditional model. Instead, it constructs the conditional model as a linear interpolation between a sampling model and the unconditional model. By training this implicit network, GFT directly optimizes the underlying sampling model, which is then employed for visual generation without guidance. In essence, one can consider GFT simply as a conditional parameterization technique in CFG training. This perspective makes GFT extremely easy to implement based on existing codebases, requiring only a few lines of modifications and with most design choices and hyperparameters inherited.
@@
Our main contribution is GFT as a fundamental visual generative modeling objective. GFT enables effective low-temperature visual generation without guided sampling. We hope it leads to a new paradigm of visual generative training.
@@ 2.1 Visual Generative Modeling
Given dataset $p({\bm{x}},{\bm{c}})$, we can train conditional diffusion models by predicting the Gaussian noise added to ${\bm{x}}_{t}$.
@@
More formally, Song et al. (2021) proved that Eq. 1 is essentially performing maximum likelihood training with evidence lower bound (ELBO). Also, the denoising model $\epsilon_{\theta}^{*}$ eventually converges to the data score function:
@@
For AR models, ${\bm{x}}_{<n}$ represents the first $i$ tokens in a pre-determined order, and ${\bm{x}}_{i}$ is the next token to be predicted. For masked models, ${\bm{x}}_{i}$ represents all the unknown tokens that are randomized masked during training, while ${\bm{x}}_{<n}$ are the unmasked ones. Due to discrete modeling, the data likelihood $p_{\theta}$ in Eq. 3 can be easily calculated.
@@ 2.2 Classifier-Free Guidance
In diffusion modeling, vanilla temperature sampling (dividing model output by a constant value) is generally found ineffective in improving generation quality (Dhariwal & Nichol, 2021). Current methods typically employ CFG (Ho & Salimans, 2022), which redefines the sampling denoising function $\epsilon_{\theta}^{\text{s}}({\bm{x}}_{t}|c)$ using two denoising models:
@@
where $\epsilon_{\theta}^{c}$ and $\epsilon_{\theta}^{\text{u}}$ respectively model the conditional data distribution $p(x|c)$ and the unconditional data distribution $p(x)$. In practice, $\epsilon_{\theta}^{\text{u}}$ can be jointly trained with $\epsilon_{\theta}^{c}$, by randomly masking the conditioning data in Eq. 1 with some fixed probability.
According to Eq. 2, CFG’s sampling distribution $p^{\text{s}}({\bm{x}}|c)$ has shifted from standard conditional distribution $p({\bm{x}}|{\bm{c}})$ to
@@
Besides diffusion, CFG is also a critical sampling technique in discrete visual modeling (Chang et al., 2022; Team, 2024; Tian et al., 2024; Xie et al., 2024).
@@ 3 Method
Despite its effectiveness, CFG requires inferencing an extra unconditional model to guide the sampling process, directly doubling the computation cost.
Moreover, CFG complicates the post-training of visual generative models because the unconditional model needs to be additionally considered in algorithm design (Meng et al., 2023; Black et al., 2023).
We propose Guidance-Free Training (GFT) as an alternative method of CFG for improving sample quality in visual generation without guided sampling. GFT matches CFG in performance but only leverages a single model $\epsilon_{\theta}^{\text{s}}({\bm{x}}|c,s)$ to represent CFG’s sampling distribution $p^{\text{s}}({\bm{x}}|c)$.
@@ 3.1 Algorithm Derivation
The key challenge in directly learning sampling model $\epsilon_{\theta}^{s}$ is the absence of a dataset that aligns with distribution $p^{\text{s}}({\bm{x}}|c)$ in Eq. 5. It is impractical to optimize maximum-likelihood-training objectives like
$\min_{\theta}\mathbb{E}_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{p^{s}({\bm{x}},{\bm{c}})}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},t,\bm{\epsilon}}\left[\|\epsilon_{\theta}^{s}({\bm{x}}_{t}|{\bm{c}})-\bm{\epsilon}\|_{2}^{2}\right].$
because we cannot draw samples from $p^{\text{s}}$.
In contrast, training $\epsilon_{\theta}^{c}({\bm{x}}|{\bm{c}})$ and $\epsilon_{\theta}^{u}({\bm{x}})$ separately as in CFG is feasible because their corresponding datasets, $\{({\bm{x}},{\bm{c}})\sim p({\bm{x}},{\bm{c}})\}$ and $\{{\bm{x}}\sim p({\bm{x}})\}$, can be easily obtained.
To address this, we reformulate Eq. 4 by simple algebra:
$\underbrace{\epsilon_{\theta}^{\text{c}}({\bm{x}}_{t}|c)}_{\begin{subarray}{c}\;\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p({\bm{x}}|{\bm{c}})\\ \text{Learnable}\end{subarray}}\;=\;\frac{1}{1+s}\underbrace{\epsilon_{\theta}^{\text{s}}({\bm{x}}_{t}|c)}_{\begin{subarray}{c}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p^{s}({\bm{x}}|{\bm{c}})\\ \text{{$\leftarrow$ \hskip 14.22636pt Target sampling model\hskip 14.22636pt +}}\end{subarray}}\;+\;\frac{s}{1+s}\underbrace{\epsilon_{\theta}^{\text{u}}({\bm{x}}_{t})}_{\begin{subarray}{c}p({\bm{x}})\\ \text{Learnable}\end{subarray}}.$ (6)
Although learning $\epsilon_{\theta}^{\text{s}}$ directly is difficult, we note it can be combined with an unconditional model $\epsilon_{\theta}^{\text{u}}$ to represent the standard conditional $\epsilon_{\theta}^{\text{c}}$, which is learnable. Thus, we can leverage the same conditional loss in Eq. 1 to train $\epsilon_{\theta}^{\text{s}}$:
@@
To this end, we have a practical algorithm for directly learning guidance-free models $\epsilon_{\theta}^{s}$. However, unlike CFG which allows controlling sampling temperature by adjusting guidance scale $s$ to trade off fidelity and diversity, our method still lacks similar inference-time flexibility as Eq. 7 is performed for a specific $s$.
To solve this problem, we define a pseudo-temperature $\beta:=1/(1+s)$ and further condition our sampling model $\epsilon_{\theta}^{s}({\bm{x}}_{t}|{\bm{c}},\beta)$ on the extra $\beta$ input. We can randomly sample $\beta\in[0,1]$ during training, corresponding to $s\in[0,+\infty)$. The GFT objective in Eq. 7 now becomes:
@@
When $\beta=1$, Eq. 8 reduces to conditional diffusion loss $\|\epsilon_{\theta}^{s}({\bm{x}}_{t}|{\bm{c}},\beta)-\bm{\epsilon}\|_{2}^{2}$. When $\beta=0$, Eq. 8 becomes an unconditional loss $\|\epsilon_{\theta}^{u}({\bm{x}}_{t})-\bm{\epsilon}\|_{2}^{2}$. This allows simultaneous training of both conditional and unconditional models.
As pseudo-temperature $\beta$ decreases $1\to 0$, the modeling target for $\epsilon_{\theta}^{s}$ gradually shifts from conditional data distribution $p({\bm{x}}|{\bm{c}})$ to lower-temperature distribution $p^{s}({\bm{x}}|{\bm{c}})$ as defined by Eq. 5.
@@ 3.2 Practical Implementation
Eq. 3.2 presents a practical loss function of GFT. The implementation is in Algorithm 1.
@@
The main difference between Eq. 3.2 and Eq. 8 is that $\epsilon_{\theta}^{u}$ is computed in evaluation mode, with model gradients stopped by the $\mathrm{\mathbf{sg}}[\cdot]$ operation. To train the model unconditionally, we randomly mask conditions ${\bm{c}}$ with $\varnothing$ when computing $\epsilon_{\theta}^{s}$. We show this design does not affect the training convergence point:
@@
Given unlimited model capacity and training data, the optimal $\epsilon_{\theta^{*}}^{s}$ for optimizing Eq. 3.2 and Eq. 8 are the same. Both satisfy
@@
Table 1: Comparison of GFT (ours) and other guidance-free methods. Numbers are reported based on experiments for the DiT-XL model or the VAR-d30 model. We use 8$\times$80GB H100 GPU cards.
@@
Loss form $\left\|\bm{\epsilon}_{\theta}^{s}-[(1\!+\!s)\bm{\epsilon}_{\phi}^{c}-s\bm{\epsilon}_{\phi}^{u}]\right\|_{2}^{2}$ $-\!\log\!\sigma(r_{\theta}^{p})\!-\!\log\!\sigma(-r_{\theta}^{n});$ $r_{\theta}\!=\!\frac{1}{s}\log\frac{p_{\theta}^{s}}{p_{\phi}^{c}}$ $\beta f_{\theta}^{s}\!+\!(1\!\!-\!\!\beta)f_{\theta}^{u}$ + Diff/AR Loss

### v2

@@ Abstract
Classifier-Free Guidance (CFG) has been a default technique in various visual generative models, yet it requires inference from both conditional and unconditional models during sampling. We propose to build visual models that are free from guided sampling. The resulting algorithm, Guidance-Free Training (GFT), matches the performance of CFG while reducing sampling to a single model, halving the computational cost. Unlike previous distillation-based approaches that rely on pretrained CFG networks, GFT enables training directly from scratch. GFT is simple to implement. It retains the same maximum likelihood objective as CFG and differs mainly in the parameterization of conditional models. Implementing GFT requires only minimal modifications to existing codebases, as most design choices and hyperparameters are directly inherited from CFG. Our extensive experiments across five distinct visual models demonstrate the effectiveness and versatility of GFT. Across domains of diffusion, autoregressive, and masked-prediction modeling, GFT consistently achieves comparable or even lower FID scores, with similar diversity-fidelity trade-offs compared with CFG baselines, all while being guidance-free. Code: https://github.com/thu-ml/GFT.
@@
Figure 2: Impact of adjusting GFT sampling temperature $\beta$ for guidance-free DiT-XL/2. GFT achieves similar results to CFG without requiring dual model inference at each step. More examples are in Figure 13.
@@ 1 Introduction
Low-temperature sampling is a critical technique for enhancing generation quality by focusing only on the model’s high-likelihood areas. Visual models mainly achieve this via Classifier-Free Guidance (CFG) (Ho & Salimans, 2022). As illustrated in Fig. 1 (left), CFG jointly optimizes the target conditional model and an extra unconditional model during training, and combines them to define the sampling process. By altering the guidance scale $s$, it can flexibly trade off image fidelity and diversity, while significantly improving the sample quality. Due to its effectiveness, CFG has been adopted as a default technique for a wide spectrum of visual generative models, including diffusion (Ho et al., 2020), autoregressive (AR) (Chen et al., 2020; Tian et al., 2024), and masked-prediction models (Chang et al., 2022; Li et al., 2023).
However, CFG is not problemless. First, the reliance on an extra unconditional model doubles the sampling cost compared with a vanilla conditional generation. Second, this reliance complicates the post-training of visual models – when distilling pretrained diffusion models (Meng et al., 2023; Luo et al., 2023; Yin et al., 2024) for fast inference or applying RLHF techniques (Black et al., 2023; Chen et al., 2024b), the extra unconditional model needs to be specially considered in the algorithm design. Third, this also renders a sharp difference from low-temperature sampling in language models (LMs), where a single model is sufficient to represent the sampling distributions across various temperatures. Similarly following LMs’ approach to divide model output by a constant temperature value is generally ineffective in visual sampling (Dhariwal & Nichol, 2021), even for visual AR models with similar architecture to LMs (Sun et al., 2024). All these lead us to ask, can we effectively control the sampling temperature for visual models using one single model?
@@
Existing attempts like distillation methods for diffusion models (Meng et al., 2023; Luo et al., 2023; Yin et al., 2024) and alignment methods for AR models (Chen et al., 2024b) are not ultimate solutions. They all rely heavily on pretrained CFG networks for loss definition and do not support training guidance-free models from scratch. Their two-stage optimization pipeline may also lead to performance loss compared with CFG, even after extensive tuning. Generalizability is also a concern. Current methods are typically tailored for either continuous diffusion models or discrete AR models, lacking the versatility to cover all domains.
@@
We propose Guidance-Free Training (GFT), a foundational algorithm for building visual generative models with no guidance. GFT matches CFG in performance while requiring only a single model for temperature-controlled sampling, effectively halving sampling costs compared with CFG. GFT offers stable and efficient training with the same convergence rate as CFG, almost no extra memory usage, and only 10–20% additional computation per training update. GFT is highly versatile, applicable in all visual domains within CFG’s scope, including diffusion, AR, and masked models.
@@
The core idea behind GFT is to transform the desired sampling model into easily learnable forms. GFT optimizes the same conditional objective as CFG. However, instead of aiming to learn an explicit conditional network, GFT defines the conditional model implicitly as the linear interpolation of a sampling network and the unconditional network (Figure 1). By training this implicit model, GFT directly optimizes the underlying sampling network, which is then employed for visual generation without guidance. In essence, one can consider GFT simply as a conditional parameterization technique in CFG training. This perspective makes GFT extremely easy to implement based on existing codebases, requiring only a few lines of modifications and with most design choices and hyperparameters inherited.
@@ 2.1 Visual Generative Modeling
Given data following $p({\bm{x}},{\bm{c}})$, we can train conditional diffusion models by predicting the Gaussian noise added to ${\bm{x}}_{t}$.
@@
More formally, Song et al. (2021) proved that Eq. (1) is essentially performing maximum likelihood training with evidence lower bound (ELBO). Also, the denoising model $\epsilon_{\theta}^{*}$ eventually converges to the data score function:
@@
For AR models, ${\bm{x}}_{<n}$ represents the first $i$ tokens in a pre-determined order, and ${\bm{x}}_{i}$ is the next token to be predicted. For masked models, ${\bm{x}}_{i}$ represents all the unknown tokens that are randomized masked during training, while ${\bm{x}}_{<n}$ are the unmasked ones. Due to discrete modeling, the data likelihood $p_{\theta}$ in Eq. (3) can be easily calculated.
@@ 2.2 Classifier-Free Guidance
Continuous CFG.
In diffusion modeling, vanilla temperature sampling (dividing model output by a constant value) is generally found ineffective in improving generation quality (Dhariwal & Nichol, 2021). Current methods typically employ CFG (Ho & Salimans, 2022), which redefines the sampling denoising function $\epsilon_{\theta}^{\text{s}}({\bm{x}}_{t}|c)$ using two models:
@@
where $\epsilon_{\theta}^{c}$ and $\epsilon_{\theta}^{\text{u}}$ respectively model the conditional data distribution $p(x|c)$ and the unconditional data distribution $p(x)$. In practice, $\epsilon_{\theta}^{\text{u}}$ can be jointly trained with $\epsilon_{\theta}^{c}$, by randomly masking the conditioning data in Eq. (1) with some fixed probability.
According to Eq. (2), CFG’s sampling distribution $p^{\text{s}}({\bm{x}}|c)$ has shifted from standard conditional distribution $p({\bm{x}}|{\bm{c}})$ to
@@
Discrete CFG.
Besides diffusion, CFG is also a critical sampling technique in discrete visual modeling (Li et al., 2023; Team, 2024; Tian et al., 2024; Xie et al., 2024). Though the guidance operation performs on the logit space instead of the score field:
$\ell_{\theta}^{\text{s}}({\bm{x}}_{n}|{\bm{x}}_{<n},{\bm{c}})=$
Given $\ell_{\theta}\propto\log p_{\theta}$, it can be proved the sampling distribution for discrete visual models also satisfies Eq. (5).
@@ 3 Method
Despite its effectiveness, CFG requires inferencing an extra unconditional model to guide the sampling process, directly doubling the computation cost. Moreover, CFG complicates the post-training of visual generative models because the unconditional model needs to be additionally considered in algorithm design (Meng et al., 2023; Black et al., 2023).
We propose Guidance-Free Training (GFT) as an alternative method of CFG for improving sample quality in visual generation without guided sampling. GFT matches CFG in performance but only leverages a single model to represent CFG’s sampling distribution $p^{\text{s}}({\bm{x}}|c)$.
@@ 3.1 Algorithm Derivation
The key challenge to directly learn the target sampling model $\epsilon_{\theta}^{s}$ is the absence of a dataset that aligns with the distribution $p^{\text{s}}({\bm{x}}|c)$ in Eq. (5). This makes it impractical to optimize a maximum-likelihood-training objective like
$\min_{\theta}\mathbb{E}_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{p^{s}({\bm{x}},{\bm{c}})}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},t,\bm{\epsilon}}\left[\|\epsilon_{\theta}^{s}({\bm{x}}_{t}|{\bm{c}})-\bm{\epsilon}\|_{2}^{2}\right],$
as we cannot draw samples from $p^{\text{s}}$. In contrast, training $\epsilon_{\theta}^{c}({\bm{x}}|{\bm{c}})$ and $\epsilon_{\theta}^{u}({\bm{x}})$ separately as in CFG is feasible because their corresponding datasets, $\{({\bm{x}},{\bm{c}})\sim p({\bm{x}},{\bm{c}})\}$ and $\{{\bm{x}}\sim p({\bm{x}})\}$, can be easily obtained.
To address this, we reformulate Eq. (4):
$\underbrace{\epsilon_{\theta}^{\text{c}}({\bm{x}}_{t}|c)}_{\begin{subarray}{c}\;\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p({\bm{x}}|{\bm{c}})\\ \text{Learnable}\end{subarray}}\;=\;\frac{1}{1+s}\underbrace{\epsilon_{\theta}^{\text{s}}({\bm{x}}_{t}|c)}_{\begin{subarray}{c}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p^{s}({\bm{x}}|{\bm{c}})\\ \text{{\hskip-28.45274pt$\leftarrow$ \hskip 14.22636pt Target sampling model\hskip 14.22636pt +\hskip-28.45274pt}}\end{subarray}}\;+\;\frac{s}{1+s}\underbrace{\epsilon_{\theta}^{\text{u}}({\bm{x}}_{t})}_{\begin{subarray}{c}p({\bm{x}})\\ \text{Learnable}\end{subarray}}.$ (6)
Although learning $\epsilon_{\theta}^{\text{s}}$ directly is difficult, we note it can be combined with an unconditional model $\epsilon_{\theta}^{\text{u}}$ to represent the standard conditional $\epsilon_{\theta}^{\text{c}}$, which is learnable. Thus, we can leverage the same conditional loss in Eq. (1) to train $\epsilon_{\theta}^{\text{s}}$, namely:
@@
To this end, we have a practical algorithm for directly learning guidance-free models $\epsilon_{\theta}^{s}$. However, unlike CFG which allows controlling sampling temperature by adjusting guidance scale $s$ to trade off fidelity and diversity, our method still lacks similar inference-time flexibility as Eq. (7) is performed for a specific $s$.
To solve this problem, we define a pseudo-temperature $\beta:=1/(1+s)$ and further condition our sampling model $\epsilon_{\theta}^{s}({\bm{x}}_{t}|{\bm{c}},\beta)$ on the extra $\beta$ input. We can randomly sample $\beta\in[0,1]$ during training, corresponding to $s\in[0,+\infty)$. The GFT objective in Eq. (7) now becomes:
@@
When $\beta=1$, Eq. (8) reduces to conditional diffusion loss $\|\epsilon_{\theta}^{s}({\bm{x}}_{t}|{\bm{c}},\beta)-\bm{\epsilon}\|_{2}^{2}$. When $\beta=0$, Eq. (8) becomes an unconditional loss $\|\epsilon_{\theta}^{u}({\bm{x}}_{t})-\bm{\epsilon}\|_{2}^{2}$. This allows simultaneous training of both conditional and unconditional models.
As pseudo-temperature $\beta$ decreases $1\to 0$, the modeling target for $\epsilon_{\theta}^{s}$ gradually shifts from conditional data distribution $p({\bm{x}}|{\bm{c}})$ to lower-temperature distribution $p^{s}({\bm{x}}|{\bm{c}})$ as defined by Eq. (5) (See Fig. 2).
@@ 3.2 Practical Implementation
A desirable algorithm should not only ensure soundness but also offer computational efficiency, seamless integration, and practical deployability. To achieve this, we further present Eq. (3.2) as a practical loss function of GFT. The implementation is in Algorithm 1.
@@
The main difference between Eq. (3.2) and Eq. (8) is that $\epsilon_{\theta}^{u}$ is computed in evaluation mode, with model gradients stopped by the $\mathrm{\mathbf{sg}}[\cdot]$ operation. To train the model unconditionally, we randomly mask conditions $\bm{c}$} with $\varnothing$ when computing $\epsilon_{\theta}^{s}$. We show this design does not affect the training convergence point:
@@
Given unlimited model capacity and training data, the optimal $\epsilon_{\theta^{*}}^{s}$ for optimizing Eq. (3.2) and Eq. (8) are the same. Both satisfy
@@
Table 1: Comparison of GFT (ours) and other guidance-free methods. Numbers are reported based on experiments for the DiT-XL model or the VAR-d30 model. We use $\times8$0GB H100 GPU cards.
@@
Loss form $\left\|\bm{\epsilon}_{\theta}^{s}-[(1\!+\!s)\bm{\epsilon}_{\phi}^{c}-s\bm{\epsilon}_{\phi}^{u}]\right\|_{2}^{2}$ $\log\!\sigma(r_{\theta}^{p})\!-\!\log\!\sigma(-r_{\theta}^{n});$ $r_{\theta}\!=\!\frac{1}{s}\log\frac{p_{\theta}^{s}}{p_{\phi}^{c}}$ $\beta f_{\theta}^{s}\!+\!(1\!\!-\!\!\beta)f_{\theta}^{u}$ + Diff/AR Loss
Reference equation Eq. (13) Eq. (14) Eq. (3.2) & Eq. (11)